A particle with a charge of -60.0 nC is placed at the center of a onconducting spherical shell of inner radius 20.0 cn and outer radius 36.0 crn. The spherical shell carries charge with a uniform density of-3.55 HC/m3. A proton moves in a circular orbit just outside the spherical shell. Calculate the speed of the proton. m/s
Answer #1

Charge at the center of spherical shell 60.0 n 7l

Charge density on spherical shell ho=-3.55,mu C/m^3

Inner radius of shell 20,0 cm CTn

Outer radius of shell b- 36.0 cm

Consider a thin spherical shell of radius r where a<r<b and thickness dr as the element.

Volume of element dV=4pi r^2dr

Charge on element dQ= ho dV=4pi ho r^2dr

Total charge on shell Q=int dQ=int ho dV=int_a^b 4pi ho r^2dr

Q=int_a^b 4pi ho r^2dr=4pi holeft.rac{r^3}{3} ight|_a^b=rac{4pi ho}{3}left(b^3-a^3 ight )

Radius of proton's orbit is b.

Apply Gauss's law to find the electric field at the location of proton.

oint_Sold{E}cdot dold{S}=rac{Q_{enc}}{epsilon_0}

Consider a Gaussian surface S of radius r where r >b. By symmetry magnitude of electric field is uniform on the surface S.Also direction of old{E} is parallel to dold{S}. Both are radially outward.

Hence oint_Sold{E}cdot dold{S}=oint_S E d{S}=Eoint_S d{S}=E(4pi r^2)

Charge enclosed Q_{enc}=q_c+Q

E=rac{Q_{enc}}{4piepsilon_0 r^2}=rac{q_c+Q}{4piepsilon_0 r^2}

When proton is just outside the shell r=b and E=rac{q_c+Q}{4piepsilon_0 b^2}

Since proton is moving in a circular orbit, rac{mv^2}{b}=e|E|

- TTno 4TEOb

v=sqrt{rac{e(|q_c|+|Q|)}{4pi epsilon_0 bm}}

v=sqrt{rac{(1.602*10^{-19})(60.0*10^{-9}+rac{4pi}{3}(3.55*10^{-6})left(36.0^3-20.0^3 ight )(10^{-6}))}{4pi (8.854*10^{-12}) (36.0*10^{-2})(1.6726*10^{-27})}}

Ξ 1.232 * 100 m/s